# Unitary discriminants of characters

### v1

@@ 3.1 Lattices
A ${\mathbb{Z}}_{L}$-lattice $\Lambda$ in $V$ is a finitely generated ${\mathbb{Z}}_{L}$ submodule of $V$ that contains a basis of $V$. The dual lattice is
@@
$\Delta(H)=\left\{\begin{array}[]{lll}(-1,-1)_{{\mathbb{Q}}}&n\equiv 2\pmod{4}&\mbox{ is ramified at }\infty,2\\ {\mathbb{Q}}&n\in 4{\mathbb{Z}}&\end{array}\right.$
@@
$\Delta(H^{\prime})=\left\{\begin{array}[]{lll}(-10,-5)_{{\mathbb{Q}}}&n\equiv 2\pmod{4}&\mbox{ is ramified at }\infty,5\\ (-10,5)_{{\mathbb{Q}}}&n\in 4{\mathbb{Z}}&\mbox{ is ramified at }2,5.\end{array}\right.$
@@ 3.2 Lattices, the local picture
4 Discriminants of characters
4.1 Characters of finite groups
Let $G$ be a finite group and $K$ be a field. A $K$-representation $\rho$ of $G$ is a group homomorphism $\rho:G\to\operatorname{GL}_{n}(K)$ into the group of invertible $n\times n$-matrices over $K$. The natural number $n$ is then called the degree of the representation. The representation $\rho$ is called irreducible, if the $KG$-module $K^{n}$ has no non-trivial $KG$-submodules. The associated Frobenius character is the map
$\chi_{\rho}:G\to K,g\mapsto\mbox{trace}(g).$
For fields $K$ of characteristic zero, Frobenius characters are in bijection to isomorphism classes of $K$-representations of $G$. As $\rho(g)$ is diagonalizable over an algebraic closure of $K$, the character values are sums of $|G|$-th roots of unity. In particular the character field ${\mathbb{Q}}(\chi_{\rho})$ of $\chi_{\rho}$, the field extension ${\mathbb{Q}}(\chi_{\rho}(g):g\in G)$ generated by the character values, is an abelian number field. The ordinary character table displays the character values of the complex irreducible characters of $G$ on representatives of the conjugacy classes of $G$. For more details see for instance [7].
To relate the representation theory over fields of characteristic 00 and positive characteristic, $p$, Brauer characters are introduced. An element of $G$ is called a $p^{\prime}$-element, if its order is not divisible by $p$. The set $G_{p^{\prime}}$ denotes the set of all $p^{\prime}$-elements in $G$; it is a union of conjugacy classes, the $p^{\prime}$-classes. A $p$-Brauer character is a certain class function $\varphi:G_{p^{\prime}}\to{\mathbb{C}}$. Loosely speaking, the restriction $\chi\pmod{p}$ of a complex Frobenius character $\chi$ to the $p^{\prime}$-classes of $G$ is a $p$-Brauer character of $G$. If $p$ does not divide the group order, then this restriction is a bijection between the irreducible $p$-Brauer characters and the irreducible complex characters. For primes $p$ dividing the group order the $p$-Brauer character $\chi\pmod{p}$ might be reducible. The irreducible constituents of $\chi\pmod{p}$ are in bijection to the composition factors of $L/\wp L$, for any lattice $L$ invariant under the representation $\rho$ affording $\chi$ and a suitably chosen prime ideal Weierstrass-$\wp$. To specify the ideal we also denote this modular reduction by $\chi\pmod{\wp}$. See [13, Sections 6 and 7.1] and [8, Introduction] for the subtleties that might arise here and Section 9 for an example. For a more sophisticated introduction to Brauer characters see [7, Chapter 15] or, for a summary, [3, Section 2.1] and [8, Introduction]. We just note the following.
@@ 4.1 Characters of finite groups
(see for instance [8, Introduction, Section 9]) Let $\chi$ be an absolutely irreducible ordinary or Brauer character of the finite group $G$ and let $\rho$ be a representation affording $\chi$. The Frobenius-Schur indicator of $\chi$ is
    if $\chi$ is real valued and there is a non-degenerate $\rho(G)$-invariant quadratic form or $\chi$ is the trivial $2$-Brauer character.
    if $\chi$ is real valued and there is a non-degenerate $\rho(G)$-invariant symplectic form but no invariant non-degenerate quadratic form.
  • o
    in all other cases.
The Frobenius-Schur indicator of an ordinary character $\chi$ is easily calculated from its character values. If the Frobenius-Schur indicator of $\chi$ is $o$ then there is a positive-definite $\rho(G)$-invariant Hermitian form for any representation $\rho$ affording $\chi$. This is not automatically true for Brauer characters, as there are irreducible Brauer characters of indicator $o$ for which the associated representation is not contained in a unitary group, for instance all indicator $o$ irreducible 3-Brauer characters of $L_{3}(3)$ are not unitary, as all irreducible $3$-modular representations of $L_{3}(3)$ are realisable over ${\mathbb{F}}_{3}$.
@@
Note that real valued Brauer characters (indicator $+$ or --) are always unitary as $\chi(g^{-1})=\overline{\chi}(g)$ so $q=p^{0}=1$ is a possible choice. The unifying philosophy here is to consider symplectic and quadratic forms as “unitary” invariant forms over a suitable field extension. Any such non-degenerate form gives rise to an isomorphism between the representation and its (unitary) dual.
Lemma 4.4.
(see [4, Lemma 4.4.1] and also [8, Introduction, Section 10] for a slightly too simplified version) An indicator $o$ absolutely irreducible Brauer character $\chi$ is unitary if and only if the representation $\rho$ affording $\chi$ admits a non-degenerate unitary form.
4.2 Discriminants of characters
Let $\chi$ be an absolutely irreducible ordinary character of a finite group $G$ of even degree $\chi(1)\in 2{\mathbb{Z}}$. Let $K:={\mathbb{Q}}(\chi)$ denote the character field of $\chi$.
If the indicator of $\chi$ is $+$ then [12] shows that there is a unique square class $\operatorname{disc}(\chi)\in K^{\times}/(K^{\times})^{2}$ such that for any extension field $L$ of $K$ and any $LG$-module $V$ affording the character $\chi$ the discriminants of all non-degenerate $G$-invariant quadratic forms $Q$ on $V$ are
$\operatorname{disc}(Q)=\operatorname{disc}(\chi)(L^{\times})^{2}.$
This square class of the character field is called the orthogonal discriminant of $\chi$ and the discriminant algebra of $\chi$ is
$\mathcal{D}(\chi):=K[X]/(X^{2}-d)\mbox{ where }\operatorname{disc}(\chi)=d(K^{\times})^{2}.$
Now assume that $\chi$ is an ordinary irreducible character of indicator $o$ and of even degree. Then the character field $L={\mathbb{Q}}(\chi)$ is a complex number field. Let $K:=L^{+}$ denote the maximal real subfield of $L$ and put $N:=N_{L/K}(L^{\times})$ to be the norm subgroup of $K^{\times}$. Then
$(K^{\times})^{2}\leq N\leq K^{\times}.$
If there is a $L$-representation $V$ of $G$ affording the character $\chi$, then there is a 1-dimensional $K$-space of $G$-invariant $L/K$-Hermitian forms spanned by some totally positive definite Hermitian form, say, $H$. In particular all non-zero $G$-invariant $L/K$-Hermitian forms have the same discriminant, $\operatorname{disc}(H)$.
@@ 4.2 Discriminants of characters
For a finite group $G$ define
$\operatorname{Irr}^{o}(G):=\{\chi\in\operatorname{Irr}_{{\mathbb{C}}}(G)\mid{\mathrm{ind}}(\chi)=o,\chi(1)\in 2{\mathbb{Z}}\}$
the set of absolutely irreducible ordinary characters of $G$ of even degree and indicator $o$.
Let $\chi\in\operatorname{Irr}^{o}(G)$, $L:={\mathbb{Q}}(\chi)$ and $K:=L^{+}$. Assume that there is an $L$-representation $V$ of $G$ affording the character $\chi$. Then the unitary discriminant of $\chi$ is defined as
$\operatorname{disc}(\chi):=\operatorname{disc}(H)=dN_{L/K}(L^{\times})\in K^{\times}/N_{L/K}(L^{\times})$
for any non-degenerate $G$-invariant Hermitian form $H$ on $V$.
The discriminant algebra of $\chi$ is defined as
$\Delta(\chi):=[(L,d)_{K}]=\Delta(H).$
As $\chi(1)$ is assumed to be even, Remark 3.4 translates into the following remark.
Remark 4.6.
If Weierstrass-$\wp$ is an infinite place of $K$, then Weierstrass-$\wp$ is ramified in $\Delta(\chi)$ if and only if $\chi(1)\equiv 2\pmod{4}$.
This accounts for the fact that $\det(\chi)$ consists of totally positive elements of $K$.
Remark 3.5 gives us
Remark 4.7.
Let $\chi\in\operatorname{Irr}^{o}(G)$. If Weierstrass-$\wp$ is a finite place of ${\mathbb{Q}}(\chi)^{+}$ that is split in the extension ${\mathbb{Q}}(\chi)/{\mathbb{Q}}(\chi)^{+}$ then Weierstrass-$\wp$ is not ramified in $\Delta(\chi)$.
Nontrivial Schur indices are very rare for indicator $o$ characters of the groups in [5] (see [6], [20]). In these rare cases we found ways to get around the direct computations by restricting the character to some appropriate subgroup. For the well-definedness of this discriminant we rely on [9, Definition (10.28), Corollary (10.35)]:
Remark 4.8.
Let ${\mathcal{A}}$ be a central simple $L$-algebra of even degree $2m$ and with involution $\iota$ of the second kind, i.e. the restriction of $\iota$ to the center $L$ of ${\mathcal{A}}$ is a field automorphism of order 2. Let $K$ denote its fixed field. Then there is a central simple $K$-algebra with involution that is Brauer equivalent to $(L,\operatorname{disc}(H))_{K}$ if $({\mathcal{A}},\iota)=(L^{2m\times 2m},\iota_{H})$ for some non-degenerate $L/K$-Hermitian form $H$ on $L^{2m}$.
4.3 Indicator --
Let $\chi$ be an absolutely irreducible complex character with Frobenius-Schur indicator --. Then the character field $K$ is a totally real number field and there is a $KG$-module $V$ affording the orthogonal character $\psi:=2\chi$. Let ${\mathcal{Q}}:=\operatorname{End}_{KG}(V)$ denote the endomorphism algebra of $V$. This is a totally definite quaternion algebra with center $K$. Choose a maximal subfield $L=K[\sqrt{-\delta}]$ of ${\mathcal{Q}}$. Then $\delta\in K$ is totally positive and the $L$-discriminant of $[{\mathcal{Q}}]$ is represented by some totally positive $\gamma\in K$ such that
${\mathcal{Q}}=(-\delta,-\gamma)_{K}.$
Remark 4.9.
Considering ${\mathcal{Q}}$ as an $L$-vector space of dimension 2, the norm form on ${\mathcal{Q}}$ is a totally positive Hermitian form of discriminant $\operatorname{disc}({\mathcal{Q}})=-\gamma$.
Proof.
Let $L=K1\oplus Ki\leq{\mathcal{Q}}$ and choose $j\in{\mathcal{Q}}$ such that $j^{2}=-\gamma$ and $ij=-ji$. Then $(1,j)$ is an orthogonal $L$-basis of ${\mathcal{Q}}$ with respect to the norm form. As the norm of $1$ is 1 and the norm of $j$ is $\gamma$ the discriminant of this 2-dimensional Hermitian $L$-space is represented by $-\gamma$. ∎
Corollary 4.10.
Let $H$ be a non-zero $G$-invariant $L/K$-Hermitian form $H$ on $V$. Then
$\operatorname{disc}(H)=\operatorname{disc}_{L}({\mathcal{Q}})^{\chi(1)/2}\mbox{ and }\Delta(H)=[{\mathcal{Q}}]^{\chi(1)/2}\in\operatorname{Br}_{2}(L,K).$
5 Elementary character theoretic methods
This section describes elementary character theoretic methods that are used to compute the discriminants of unitary ordinary characters of $G$.
5.1 Unitary stability
Definition 5.1.
Similarly as in [13, Theorem 5.13] one has the following proposition.
Proposition 5.2.
An ordinary character is unitary stable, if and only if it has a well defined unitary discriminant algebra.
The unitary discriminant of a sum of unitary stable characters is essentially the product of the unitary discriminants of the unitary summands. If these summands have a larger character field, then also their Galois conjugates appear and one needs to replace the product by the appropriate norm as in [13, Proposition 5.17], Summands of indicator -- are handled in Corollary 4.10. Similarly Schur indices for summands $\chi$ of indicator $+$ give rise to an additional correction factor, $\operatorname{disc}_{L}({\mathcal{Q}})^{\chi(1)/2}$ (see Theorem 7.4).
Remark 5.3.
In our computations we often get equations for discriminants of unitary stable characters that hold over the complex field $L$. This means that the equations hold in $K^{\times}/N_{L/K}(L^{\times})$ where $K=L^{+}$ is the totally real subfield of $L$. If there is no field mentioned we always mean the character field.
5.2 Modular reduction
Theorem 5.4.
Let $\chi\in\operatorname{Irr}^{o}(G)$ and put $L:={\mathbb{Q}}(\chi)$. Let Weierstrass-$\wp$ be a finite place of $L^{+}$ that is inert in $L/L^{+}$. If $\chi\pmod{\wp}$ is unitary stable then Weierstrass-$\wp$ is not ramified in the discriminant algebra $\Delta(\chi)$.
Proof.
The proof of this theorem is almost the same as the one of [13, Theorem 6.4], showing the result for real characters $\chi$. Put $K=L^{+}$ to denote the real subfield of $L$. As Weierstrass-$\wp$ is inert, the completion $L_{\wp}$ of $L$ at the finite place Weierstrass-$\wp$ of $K$ is the unique unramified extension of degree $2$ of $K_{\wp}$. Denote by $P$ the maximal ideal of ${\mathbb{Z}}_{L_{\wp}}$ and by $F$ the residue field ${\mathbb{Z}}_{L_{\wp}}/P$. After possibly passing to a suitable unramified extension of $K_{\wp}$ we may assume that there is an $L_{\wp}G$-module $(V,H)$ affording the character $\chi$. Let $\Lambda\subseteq\Lambda^{*}$ be a maximal integral $G$-invariant lattice in $V$. Then
$P\Lambda^{*}\subseteq\Lambda\subseteq\Lambda^{*}$
and $(\Lambda^{*}/\Lambda,\tilde{H})$ is a Hermitian $FG$-module by Proposition 3.10 (e). Now $\Lambda$ is maximal integral, so this $FG$-module is the orthogonal direct sum of simple unitary modules. By assumption all these modules have even dimension, so $\dim_{F}(\Lambda^{*}/\Lambda)$ is even and Proposition 3.10 (d) shows that the Weierstrass-$\wp$-adic valuation of the discriminant of $H$ is even. ∎
Corollary 5.5.
Let $\chi\in\operatorname{Irr}^{o}(G)$. Then all prime ideals of ${\mathbb{Q}}(\chi)^{+}$ that ramify in the discriminant algebra $\Delta(\chi)$ divide $2|G|$.
For primes that are ramified in $L/L^{+}$ the module $\Lambda^{*}/\Lambda$ carries a symplectic $G$-invariant form. So Proposition 3.10 (c) implies the following result.
Corollary 5.6.
Let $\chi\in\operatorname{Irr}^{o}(G)$. If a non-dyadic prime Weierstrass-$\wp$ is ramified in ${\mathbb{Q}}(\chi)/{\mathbb{Q}}(\chi)^{+}$ and $\chi\pmod{\wp}$ is unitary stable, then $\operatorname{disc}(\chi)\pmod{\wp}$ is the orthogonal discriminant of $\chi\pmod{\wp}$. This implies that Weierstrass-$\wp$ is ramified in the discriminant algebra $\Delta(\chi)$ if and only if the orthogonal discriminant of $\chi\pmod{\wp}$ is not a square in the residue field.
For blocks of defect $1$, also the converse of Theorem 5.4 is true. For general cyclic defect an analogous statement as [13, Theorem 6.12] is also true for unitary stable characters.
Proposition 5.7.
Let $\chi\in\operatorname{Irr}^{o}(G)$ be in a $p$-block of defect 1. Put $L:={\mathbb{Q}}(\chi)$ and let Weierstrass-$\wp$ be a finite place of $L^{+}$ that contains the rational prime $p$ and is inert in $L/L^{+}$. Then $\chi\pmod{\wp}$ is unitary stable if and only if Weierstrass-$\wp$ does not ramify in the discriminant algebra of $\chi$.
The proof is almost the same as for orthogonally stable characters in [13, Theorem 6.10]. In particular this shows that if Weierstrass-$\wp$ is inert and $\chi$ in a block of defect 1, then the reduction $\chi\pmod{\wp}$ has at most two unitary constituents.
Example 5.8.
As an example let $\chi$ be one of the two complex conjugate indicator $o$ irreducible ordinary faithful character of degree $116622$ of the group $3.ON$, the Schur cover of the sporadic simple O’Nan group of order $2^{9}\cdot 3^{5}\cdot 5\cdot 7^{3}\cdot 11\cdot 19\cdot 31$. Then ${\mathbb{Q}}(\chi)={\mathbb{Q}}[\sqrt{-3}]$ and $7,19$ and $31$ are norms in ${\mathbb{Q}}[\sqrt{-3}]/{\mathbb{Q}}$. The 11-modular reduction of $\chi$ is irreducible. Modulo $5$ the character has defect $1$ and $\chi\pmod{5}=5643ab+52668ab$. Whereas $5643a$ and $5643b$ are unitary, the other two 5-modular constituents do not admit a non-zero invariant form. So Proposition 5.7 implies that $5$ divides the discriminant of $\chi$ and hence $\operatorname{disc}(\chi)\in\{-5,-10\}$. Also the 3-modular reduction $\chi\pmod{3}=6138+104346$ is unitary stable. Both 3-modular constituents have indicator $+$, $\operatorname{disc}(6138)=O-$ and $\operatorname{disc}(104346)=O+$. So the unitary discriminant of $\chi$ is not a square modulo $3$, and hence $\operatorname{disc}(\chi)=-10$.
5.3 The unitary discriminants for $O_{10}^{+}(2)$.
Theorem 5.9.
$\begin{array}[]{|c|c|c|c|c|}\hline\cr\chi&\chi(1)&{\mathbb{Q}}(\chi)&\operatorname{disc}(\chi)&\Delta(\chi)\\ \hline\cr 33,34&110670&{\mathbb{Q}}[\sqrt{-15}]&-1&[(-1,-3)_{{\mathbb{Q}}}]\\ \hline\cr 51,52&332010&{\mathbb{Q}}[\sqrt{-15}]&-2&[(-2,-5)_{{\mathbb{Q}}}]\\ \hline\cr 68,69&442680&{\mathbb{Q}}[\sqrt{-15}]&1&[{\mathbb{Q}}]\\ \hline\cr 79,80&711450&{\mathbb{Q}}[\sqrt{-7}]&-3&[(-1,-3)_{{\mathbb{Q}}}]\\ \hline\cr 81,82&711450&{\mathbb{Q}}[\sqrt{-7}]&-3&[(-1,-3)_{{\mathbb{Q}}}]\\ \hline\cr\end{array}$
Proof.
The first three pairs of characters $\chi$ with character field $L:={\mathbb{Q}}[\sqrt{-15}]$ have a unitary stable reduction modulo all primes that are inert in $L/{\mathbb{Q}}$. So by Theorem 5.4 no inert prime ramifies in the discriminant algebra of $\chi$. Remark 4.7 tells us that no split prime ramifies in $\Delta(\chi)$ and Remark 4.6 shows that the infinite place of ${\mathbb{Q}}$ ramifies in $\Delta(\chi)$ for the first two but not for the third of the three pairs of characters. For all three characters the 5-modular reduction is absolutely irreducible. From the database of orthogonal discriminants in [1] we get that this 5-modular reduction has orthogonal discriminant $O+$ for $\chi_{33/34}$ and $\chi_{68/69}$ and $O-$ for $\chi_{51/52}$. So Corollary 5.6 shows that $5$ is ramified in $\Delta(\chi)$ only for $\chi=\chi_{51/52}$. As the number of ramified places in $\Delta(\chi)$ is even, this yields the behaviour of the prime $3$. So $\Delta(\chi_{33})$ is ramified at $\infty$ and $3$, $\Delta(\chi_{51})$ is ramified at $\infty$ and $5$, and $\Delta(\chi_{61})=[{\mathbb{Q}}]$.
The four characters of degree 711450 are absolutely irreducible modulo $5,7,17,$ and $31$. Modulo the ramified prime, 7, their orthogonal discriminant is $O+$ according to [1]. As 2 is a norm in ${\mathbb{Q}}[\sqrt{-7}]/{\mathbb{Q}}$ the possible discriminants are $-1$ and $-3$. But $-1$ is not a square modulo $7$ and hence $\operatorname{disc}(\chi)=-3$ for these four characters $\chi$. ∎
5.4 Restriction and induction
Often the degree of $\chi$ is too big so that we cannot construct such a representation $\rho$, however sometimes the restriction of $\chi$ to a subgroup remains unitary stable and all its summands that occur in $\chi_{|U}$ with odd multiplicity have manageable degree.
Remark 5.10.
If there is a subgroup $U\leq G$ such that the restriction $\chi_{|U}$ of $\chi$ to $U$ is unitary stable, then $\operatorname{disc}(\chi)=\operatorname{disc}(\chi_{|U})$.
Note that the parity of the degrees of the unitary constituents of the restriction matters, but their character fields are not relevant. For induction it is the opposite: If $\psi$ is a character of some subgroup $U$ and $F$ a $U$-invariant form in a representation affording the character $\psi$, then the orthogonal sum of $[G:U]$ copies of $F$ is a $G$-invariant form in the induced representation.
Remark 5.11.
Let $U\leq G$ and let $\psi$ be some character of $U$. If the induced character $\psi^{G}$ is unitary stable then
$\operatorname{disc}(\psi^{G})=\left\{\begin{array}[]{ll}\operatorname{disc}(\psi)&\mbox{ if }[G:U]\mbox{ odd }\\ 1&\mbox{ if }[G:U]\mbox{ even }\end{array}\right.$
This equation holds over the character field of $\psi$, which might be larger than the one of $\psi^{G}$.
5.5 Tensor products and symmetrizations
Remark 5.12.
Let $\chi,\psi$ be two absolutely irreducible ordinary characters of $G$ such that $\chi(1)$ is even. If the tensor product $\chi\cdot\psi$ is unitary stable, then
$\operatorname{disc}(\chi\cdot\psi)=\operatorname{disc}(\chi)^{\psi(1)}$
over any field containing ${\mathbb{Q}}(\chi)$ and ${\mathbb{Q}}(\psi)$.
Symmetrizations have been dealt with in [14] and illustrated with the example of the group $6.Suz$, where the symmetrizations of the complex character of degree 12 allow to conclude that all faithful $\chi\in\operatorname{Irr}^{o}(6.Suz)$ have unitary discriminant $(-1)^{\chi(1)/2}$ (see [14, Section 6]). Symmetrizations are quite helpful for groups $G$ for which one of its covering groups has a faithful character $\chi$ of fairly small degree. As for induced characters it is not important that $\chi(1)$ is even to use unitary stable symmetrizations of $\chi$ for getting equations for unitary discriminants over the character field of $\chi$.
6 Perfect groups with even center
Proposition 6.1.
Let Weierstrass-$\wp$ be a non-dyadic prime of $K$.
@@ 6 Perfect groups with even center
    If Weierstrass-$\wp$ is unramified in $L/K$ then Weierstrass-$\wp$ is unramified in $\Delta(\chi)$.

### v2

@@ 3.1 Lattices
A ${\mathbb{Z}}_{L}$-lattice $\Lambda$ in $V$ is a finitely generated ${\mathbb{Z}}_{L}$-submodule of $V$ that contains a basis of $V$. The dual lattice is
@@
[…]
@@
[…]
@@ 3.2 Lattices, the local picture
4 Algebras with involution
I thank the referee for the motivation to set up a more sophisticated theory for the definition of discriminant algebras.
4.1 Central simple algebras with involution
This section briefly recalls some notations for algebras with involution over fields of characteristic $\neq 2$. A much more sophisticated treatment can be found in [10].
Definition 4.1.
Let $L$ be a field of characteristic $\neq 2$ and let $(A,\iota)$ be a central simple $L$-algebra with an involution $\iota$, i.e. a ring anti-automorphism of order $2$. Then $\dim_{L}(A)=n^{2}$ is a square.
  • (i)
    The degree of $A$ is $\deg(A):=n$.
  • (ii)
    The involution is said to be of the first kind, if $\iota_{|L}=\operatorname{id}_{L}$ and of the second kind, if the restriction to the center $L$ is non-trivial.
  • (iii)
    Put $A^{+}:=\{a\in A\mid\iota(a)=a\}$ and $A^{-}:=\{a\in A\mid\iota(a)=-a\}$. Then $A=A^{+}\oplus A^{-}$.
  • (iv)
    An involution of the first kind is called orthogonal, if $\dim_{L}(A^{-})=n(n-1)/2$ and symplectic, if $\dim_{L}(A^{-})=n(n+1)/2$.
The involutions of matrix rings $A=L^{n\times n}$ are all adjoint involutions of Hermitian, symmetric or skew-symmetric forms, and these are of the second kind, respectively orthogonal or symplectic.
For any central simple $L$-algebra AA there is a field extension $E$, such that $E\otimes_{L}A\cong E^{n\times n}$. For $a\in A\subseteq E\otimes_{L}A\cong E^{n\times n}$ we define the reduced norm of $a$ as the determinant of the matrix $a\in E^{n\times n}$, $N_{red}(a)=\det(a)$ (see [18, Section 9]). This definition depends neither on the chosen splitting field $E$ nor on the identification with a matrix ring. Moreover for any $a\in A$ we have $N_{red}(a)\in L$.
Discriminants of involutions are only well defined if the degree of $A$ is even:
@@ 4.1 Central simple algebras with involution
([10, Definition 7.2], [9]) $A$ be a central simple $L$-algebra of even degree $2n$ and $\iota$ be an orthogonal involution on $A$. Then $A^{-}$ contains invertible elements. The discriminant of $\iota$ is defined as the square class
$\operatorname{disc}(\iota):=(-1)^{n}N_{red}(a)(L^{\times})^{2}$
for any invertible $a\in A^{-}$.
In the split case, $A=L^{2n\times 2n}$ and a symmetric $B\in\operatorname{GL}_{2n}(K)$ such that
$\iota=\iota_{B},X\mapsto BX^{tr}B^{-1}$
is the adjoint involution with respect to $B$, we have
Similar to the orthogonal case, there is a well defined discriminant algebra for even degree algebras with unitary involutions:
@@
([10, Definition (10.28), Corollary (10.35)]) Let $\mathcal{A}$} be a central simple $L$-algebra of even degree $2n$ and with an involution $\iota$ of the second kind Let $K:=\operatorname{Fix}_{L}(\iota_{|L})$ denote the fixed field of $\iota$ in $L$. Then we put the discriminant algebra $\Delta(({\mathcal{A}},\iota))]\in\operatorname{Br}(K)$ of $({\mathcal{A}},\iota)$ to denote the Brauer class of the central simple $K$-algebra $\Delta(({\mathcal{A}},\iota))$ with involution described in [10, Section 10].
The discriminant algebra $\Delta(({\mathcal{A}},\iota))$ is constructed in [10, Section 10] as a $K$-subalgebra of the central simple $L$-algebra $\lambda^{n}{\mathcal{A}}=L\otimes_{K}\Delta(({\mathcal{A}},\iota))$. The latter is Brauer equivalent to ${\mathcal{A}}]^{n}\in\operatorname{Br}(L)$. This algebra is split, ${\mathcal{A}}]^{n}=[L]\in\operatorname{Br}(L)$, if and only if $L$ splits $\Delta(({\mathcal{A}},\iota))$. For number fields $L$ this implies that $L$ is a maximal subfield of the quaternion algebra in $\Delta(({\mathcal{A}},\iota))]$. Then the discriminant of $({\mathcal{A}},\iota)$ is defined as the $L$-discriminant of $\Delta(({\mathcal{A}},\iota))]$.
@@
Assume that $L$ is a number field. A central simple $L$-algebra $\mathcal{A}$} of degree $2n$ is said to be quasi-split, if ${\mathcal{A}}]^{n}=[L]\in\operatorname{Br}(L)$.
For an $L/K$-Hermitian involution $\iota$ on such a quasi-split $L$-algebra $\mathcal{A}$} we define the unitary discriminant
$\operatorname{disc}(({\mathcal{A}},\iota)):=\operatorname{disc}_{L}([\Delta(({\mathcal{A}},\iota))])=\delta N_{L/K}(L^{\times})$
if $\Delta(({\mathcal{A}},\iota))]=[(L,\delta)_{K}].$
I thank J.P. Tignol for communicating to me that [11, Corollary 2.12] shows that there is an extension $F$ of $K$ splitting $\mathcal{A}$} and such that the map $\operatorname{Br}(K)\to\operatorname{Br}(F),[D]\mapsto[D\otimes_{K}F]$ is injective. This observation is also used in the proof of [10, Corollary (10.36)]. It allows to conclude the following remark.
Remark 4.5.
The discriminant algebra $\Delta(({\mathcal{A}},\iota))]$ is the unique element of $\operatorname{Br}(K)$ with the following property: For any extension $E^{+}$ of $K$ such that $E:=E^{+}L$ is a splitting field of $\mathcal{A}$} and a Hermitian form $H$ on $E^{2n}$ with $(\iota_{H})_{|{\mathcal{A}}}=\iota$ we get
$E^{+}\otimes_{K}\Delta(({\mathcal{A}},\iota))]==\Delta(H)=[(E,\operatorname{disc}(H))_{E^{+}}]\in\operatorname{Br}(E^{+}).$
Note that Remark 4.5 characterises the unitary discriminant of a quasi-split central simple $L$-algebra $({\mathcal{A}},\iota)$ by the following property: For any extension $E^{+}$ of $K$ such that $E:=E^{+}L$ is a splitting field of $\mathcal{A}$} and a Hermitian form $H$ on $E^{2n}$ with $(\iota_{H})_{|{\mathcal{A}}}=\iota$ we get ${\mathcal{A}}\otimes E=E^{2n\times 2n}$ and
$\operatorname{disc}(({\mathcal{A}},\iota))N_{E/E^{+}}(E^{\times})=\operatorname{disc}(H).$
4.2 Orthogonal subalgebras
@@
Let $\mathcal{A}$} be a central simple $L$-algebra of even degree and with $L/K$-involution $\iota$ of second kind. Then a $K$-subalgebra AA of $\mathcal{A}$} is called an orthogonal subalgebra of $({\mathcal{A}},\iota)$, if and only if
@@
    AA is a central simple $K$-algebra with $LA={\mathcal{A}}$.